Taking off from a point on the equator in the late afternoon and flying due west, passengers on the concorde supersonic airliner see the sun set and then rise again ahead of them. Estimate concorde’s minimum speed. (Earth’s radius=6400km)
1
Expert's answer
2019-07-15T09:25:24-0400

the velocity must be greater than the linear velocity of a point at the equator, the minimum value

"v = \omega R = \frac{2\pi \cdot6.4}{24\cdot3.6} = 0.47 (km/s)"

or 1700 km/h